# In-context Prompt Learning for Test-time Vision Recognition with Frozen Vision-language Model

### v1

@@ 4.3 Ablation Studies
Component Analysis. In this experiment, we assess the impact of each component within our method on both task and instance adaptation. With only a limited number of in-context examples, task-specific adaptation employs them as training data to fine-tune a task-aware prompt, which can then be applied to any test sample from a new task. In contrast, sample-specific adaptation does not depend on in-context examples; instead, it learns a sample-aware prompt using only unlabeled test samples. To investigate the effect of task- and instance-specific adaptation, We create three distinct variants of our method: (1) Ours without supervised objective (w/ U-obj) does not adapt task-specific labeled examples for each task in prompt learning; (2) Ours without unsupervised objective (w/o U-obj) does not utilize on-the-fly unlabeled test sample to learn sample-specific patterns; (3) Ours with context-aware unsupervised objective (CU-obj) adopt in-context examples as domain-specific context information and employ test sample to query model to directly predict the output label.
Results are reported in Table 4. The two methods without supervised/unsupe-rvised objectives show lower performance than ours with context-aware objective. This demonstrates that incorporating task-specific and instance-specific adaptation are both beneficial for fine-grained vision recognition problems. Task-specific adaptation is a crucial component of our problem, as instance-specific adaptation is limited to tuning prompt tokens solely on unlabeled samples using unsupervised objectives. Our method learns prompt tokens directly from task-specific image-label pairs and instance-specific test sample to adapt model accordingly.
Effect on In-context Examples and Unlabeled Test Sample. In-context examples do not belong to the same categories as the unlabeled test samples. Consequently, they provide task-specific cross-modality alignment information rather than class-specific information, as is typical in traditional few-shot learning. The test samples are unlabeled and only contribute visual modality information to the CLIP model, lacking any language information.
As shown in Table 5, visual prompt learning significantly improves the zero-shot performance of CLIP compared to language prompt learning. We also evaluate the result of our InCPL when visual and language prompts are concurrently tuned, referred to as “concurrent-prompt”, which yields inferior performance compared to single visual prompt learning. Further, we consider an image-label pair from the same categories with test sample and conduct one-shot learning
@@
InCPL w/ V-P 1-shot 71.99 46.57 90.35
InCPL w/ T-P 1-shot 76.65 50.77 90.81
InCPL w/ Con-P 1-shot 76.09 51.60 92.04
InCPL w/ Cyc-P 1-shot 82.62 53.84 95.86
for our approach. The one-shot can provide class-specific cross-modality alignment information for unlabeled test sample. Compared to in-context learning, one-shot learning significantly improves the performance of InCPL w/ concurrent-prompt. We argue that concurrent-prompt does not explore the class-specific contextual information of one-shot sample. As shown in Table 5, the proposed cyclic prompt learning achieve the best accuracies and outperform concurrent-prompt with large margin (e.g., +6.53 for Flower dataset). This can be attributed to the disentanglement of contextual information into different modalities through cyclic learning. Here we perform a single iteration of visual and language prompt with sequential learning, and additional iterations are expected to yield improved results. We leave the exploration of iterative training for future research.
Table 6: Accuracy of different visual prompt methods.
Method Flower DTD Pets
Top 1 acc. Top 1 acc. Top 1 acc.
Patched prompt [2] 57.90 32.51 77.98
Padded prompt [15] 56.07 32.68 79.39
Token prompt 59.32 34.63 74.19
Generic-language prompt 63.30 44.74 85.20
Ours (Language-aware) 71.13 47.34 90.25
@@
Comparison with Previous Visual Prompt Approaches. Apart from our proposed language-aware visual prompting, we also evaluate different visual prompting methods, including pixel-level patched, padded prompts [2, 15] and sequence-level token prompt with random initialization. As shown in Figure 6, we learn a single image perturbation by adding learnable patched and padded prompt to input image, which are respectively denoted as “patched prompt” and “padded prompt”. Following [15], we also prepend some learnable parameters into input sequence of transformer, denoted as “token prompt”. Table 6 shows the accuracies with different prompt designs. Compared with patched-prompt, padded-prompt, and token-prompt, our proposed visual prompt provides a robust language-aware initialization for model optimization, which not only enhances prompt-based performance but also contributes to increased training stability. To further examine the impact of parameter initialization, we experiment with generic-language prompt, which uses a generic language initialized with random context vectors. As shown in Table 6, generic-language prompt show inferior performance to our language-aware prompt but outperforms the language-unaware prompt methods (i.e., padded, patched and token prompts). This demonstrates that incorporating language modality information into visual branch is beneficial to prompt-based visual recognition.
@@
Does the number of in-context examples matter? Recall that in-context examples are sampled from the demonstration pool, which contains one sample from each category. We are interested in understanding whether the quantity of these examples impacts performance, particularly for the supervised prompt retrieval method. To investigate this, we varied the number of in-context examples from 1 to 19, resulting in a set of results illustrated in Figure 5. Firstly, it’s evident that using in-context examples significantly outperforms the “no examples” method, and the performance clearly benefits from a larger number of examples. However, interestingly, increasing the number of examples beyond a certain point (i.e., 5) starts to decrease their performance. This phenomenon may be attributed to the fact that in-context examples can be considered as a form of “training data” , and an abundance of training data may lead the model to learn prompts specific to the in-context examples rather than the unlabeled test samples. While having more examples may not necessarily be advantageous, the crucial question lies in how to strategically select and employ these examples to harness the full potential of the CLIP model. For a comprehensive exploration and discussion of this topic, please refer to the Section 6.2 in the Appendix.
Ground Truth Matters. To study the impact of correctly-paired inputs and labels in the in-context examples, referred as “ground truth input-label mapping”, we evaluate the following three methods.
• No example is a typical test-time adaptation method that does not use any labeled data. A prediction is made via optimizing unsupervised objective involving unlabeled test sample.
• Examples w/o labels is the baseline that only uses input images without the corresponding labels, which indicates the model performance without looking at the label information.
• Examples w/ gold labels are employed in a typical in-context learning method with a set of labeled examples, which indicates the model performance with looking at relevant knowledge.
• Examples w/ random labels replace all gold labels with random labels, which are randomly sampled at uniform from label space on the testing data.
@@ 5 Concusion
Limitation. We develop a novel in-context evaluation paradigm that improve the performance of existing prompt learning algorithms in a ChatGPT-like manner. Under this paradigm, the in-context evaluator operates on test samples at test-time with unaltered model parameters, where the model only process unlabeled test sample and in-context examples served as context information, without any other samples from the same category. The proposed InCPL requires a task-specific context dataset for each individual task. This dataset creation process inherently introduces certain priors, which can potentially limit the broad applicability of our proposed methods. In future work, we will consider developing unique context datasets tailored to different tasks. This approach would enable the adaptive selection of image-label pairs to serve as in-context prompt information for various tasks.

### v2

@@ 4.4 Ablation Studies
Component Analysis. In this experiment, we assess the impact of each component within our method on both task and instance adaptation. With only a limited number of in-context examples, task-specific adaptation employs them as training data to fine-tune a task-aware prompt, which can then be applied to any test sample from a new task. In contrast, sample-specific adaptation does not depend on in-context examples; instead, it learns a sample-aware prompt using only unlabeled test samples. To investigate the effect of task- and sample-specific adaptation, We create three distinct variants of our method: (1) Ours with unsupervised objective (w/ U-obj) utilize on-the-fly unlabeled test sample to learn sample-specific patterns; (2) Ours without supervised objective (w/ S-obj) adopts task-specific labeled examples for each task in prompt learning; (3) Ours with context-aware unsupervised objective (CU-obj) adopts labeled examples as task-specific context information and employs test sample to query the prior knowledge of pre-trained model to predict its label.
All the ablation studies are conducted over fined-grained datasets and OOD ImageNet datasets. As shown in Table 4, our InCPL (w/ CU-obj) significantly surpass other baselines, demonstrating that test-time adaptation can be improved by introducing task-specific context information with labeled examples. Besides, task- and sample-specific designs in InCPL can complement each other as InCPL (i.e., the combination of the two designs) clearly outperforms either task-specific adaptation or sample-specific adaptation on the challenging fine-grained dataset. Moreover, we extend our ablation studies to the OOD benchmark, and results on ImageNet-R dataset show that the task-specific adaptation achieved 76.86% accuracy, the sample-specific adaptation 75.73%, while their combination yielded a 77.56% accuracy, thereby highlighting the significance of each kind of adaptation in enhancing InCPL’s performance.
Table 5: Accuracy of different visual prompt methods.
Method Flower DTD Pets Imagenet-R ImageNet-S
Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc.
Patched prompt 57.90 32.51 77.89 70.62 43.69
Padded prompt 56.07 32.68 79.39 68.54 40.48
Token prompt 59.32 34.63 74.19 70.39 40.1
Generic-language prompt 63.30 44.74 85.2 76.46 45.29
Ours (Language-aware) 71.13 47.34 90.25 77.56 48.03
@@
InCPL w/ V-P One-shot 71.99 46.57 90.35
InCPL w/ T-P One-shot 76.65 50.77 90.81
InCPL w/ Con-P One-shot 76.09 51.60 92.04
InCPL w/ Cyc-P One-shot 82.62 53.84 95.86
@@
Comparison with Previous Visual Prompt Approaches. Apart from our proposed language-aware visual prompting, we also evaluate different visual prompting methods, including pixel-level patched, padded prompts [1, 15] and token prompt with random initialization. As shown in Figure 5, we learn a single image perturbation by adding learnable patched and padded prompt to input image, which are respectively denoted as “patched prompt” and “padded prompt”. Following [15], we also prepend some learnable parameters into input sequence of transformer, denoted as “token prompt”. Table 5 shows the accuracies with different prompt designs. Compared with patched-prompt, padded-prompt, and token-prompt, our proposed visual prompt provides a robust language-aware prompt initialization for model optimization, which not only enhances prompt-based performance but also contributes to increase training stability. To further examine the impact of parameter initialization, we experiment with generic-language prompt, which is randomly initialized with a generic context. As shown in Table 5, generic-language prompt show inferior performance to our language-aware prompt but outperforms the language-unaware prompt methods (i.e., padded, patched and token prompts). This demonstrates that incorporating language modality information into visual branch is beneficial to prompt-based visual recognition.
Comparison with Different Prompt Learning Strategies under In-context Examples and One-shot Scenarios. In our InCPL, in-context examples do not belong to the same categories as the unlabeled test samples. Consequently, they provide task-specific cross-modality alignment information rather than class-specific information, as is typical in traditional few-shot learning. As a comparison, we also use one-shot samples to provide class-specific information during test-time prompt learning.
As shown in Table 6, visual prompt learning significantly improves the zero-shot performance of CLIP compared to language prompt learning. We also evaluate the result of our InCPL when visual and language prompts are concurrently tuned, referred to as “concurrent-prompt”, which yields inferior performance compared to single visual prompt learning. We argue that this is mainly because in-context examples only provide context information and do not offer cross-modal alignment information about the test samples. To further investigate this, we further consider an image-label pair from the same categories with test sample and conduct one-shot learning for our approach. The one-shot can provide class-specific cross-modality alignment information for unlabeled test sample. Compared to in-context learning, one-shot learning significantly improves the performance of InCPL w/ concurrent-prompt. As shown in Table 6, the proposed cyclic prompt learning achieve the best performance and outperform concurrent-prompt with large margin (e.g., +6.53 for Flower dataset). This can be attributed to the disentanglement of contextual information into different modalities through cyclic learning, enabling the model to better utilize labeled data. Here we perform a single iteration of visual and language prompt with sequential learning, and more iterations are expected to yield improved results. We leave the exploration of iterative training for future research.
@@
Does the Number of In-context Examples Matter? Recall that in-context examples are sampled from the demonstration pool, which contains one sample from each category. We are interested in understanding whether the quantity of these examples has important impacts on model performance. To investigate this, we varied the number of in-context examples from 1 to 19, resulting in a set of results illustrated in Figure 5. Firstly, it is evident that using in-context examples significantly outperforms the ”no examples” method, and the performance clearly benefits from a larger number of examples. However, interestingly, increasing the number of examples beyond a certain point (i.e., 5) starts to decrease performance. This phenomenon may be attributed to the fact that in-context examples can be considered as a form of “training data,” and an abundance of training data may lead the model to learn prompts specific to the in-context examples rather than the unlabeled test samples. While having more examples may not necessarily be advantageous, the crucial question lies in how to strategically select and employ these examples to harness the full potential of the CLIP model. For a comprehensive exploration and discussion of this topic, please refer to Section 4.5.
Ground Truth Matters. To study the impact of correctly-paired inputs and labels in the in-context examples, referred to as “ground truth input-label mapping”, we evaluate the following three methods.
• No example is a typical test-time adaptation method that does not use any labeled data. A prediction is made by optimizing an unsupervised objective involving the unlabeled test sample.
• Examples w/o labels is the baseline that only uses input images without the corresponding labels, indicating the model’s performance without considering label information.
• Examples w/ gold labels are employed in a typical in-context learning method with a set of labeled examples, indicating the model’s performance with access to relevant knowledge.
• Examples w/ random labels replace all gold labels with random labels, which are randomly sampled uniformly from the label space of the testing data.
@@
Figure 7: Comparison with few shot learning when using few in-context examples on Flower102.
Table 7: Comparison with CoOp and CoCoOp using the same examples.
Method Flower102 DTD Pets Cars Caltech101
Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc.
CoOp [50] 66.10 30.97 82.77 60.20 90.26
CoCoOP [49] 67.23 31.72 83.14 59.78 90.43
Ours 71.13 47.34 90.60 67.54 94.69
Figure 8: Experimental results w.r.t. different in-context example selection strategies.
Table 8: Comparative analysis of inference time and accuracy with existing TPT. The inference time (Infer. Time) is calculated in minutes.
Method Flower102 Pets Cars Caltech101
Infer. Time ()($\downarrow$)( ↓ ) Top 1 acc. ()($\uparrow$)( ↑ ) Infer. Time ()($\downarrow$)( ↓ ) Top 1 acc. ()($\uparrow$)( ↑ ) Infer. Time ()($\downarrow$)( ↓ ) Top 1 acc. ()($\uparrow$)( ↑ ) Infer. Time ()($\downarrow$)( ↓ ) Top 1 acc. ()($\uparrow$)( ↑ )
TPT [32] 97.22 68.98 111.31 87.79 322.59 66.87 86.25 94.16
InCPL (Ours) 23.79 72.27 16.31 90.62 69.05 67.54 35.54 94.69
4.5 Additional Ablation Studies
Comparison with CoOP and CoCoOP using the same examples. We provide CoOp/CoCoOp’s results using the same examples as InCPL in Table 7. Experimental results show that our InCPL achieves better performance than CoOp and CoCoOp on this setting.
Training Strategy: In-context Learning vs. Few-shot Learning. A conventional approach for leveraging the in-context examples dataset involves fine-tuning the prompt on a labeled dataset. As an alternative method, we implemented few-shot prompt learning using all available in-context examples and conducted a comparison with our InCPL approach. As demonstrated in Figure 7, the image representations learned through the few-shot approach exhibit lower resolution in distinguishing between classes compared to our prompted representations. The few-shot approach might introduce more variance and noise into the feature space, and the learned features are specific to the task rather than the individual test sample. In contrast, our InCPL method is specifically designed to learn an informative prompt for each test sample using in-context examples, enabling a more effective alignment between the sample and its associated examples.
In-context Example Selection: Random vs. Definition. This experiment aims to evaluate two distinct approaches for selecting context examples to prompt each test sample: random and definition-based approaches. In the former approach, input-label pairs are randomly selected from candidate examples as in-context examples for each test sample, while the latter approach utilizes a common set of examples shared across all test samples once in-context examples have been sampled. As depicted in Figure  8, the definition-based selection approach also exhibits significant fluctuations with different random seeds, and its performance consistently lags behind the random-based selection approach. The primary reason for this discrepancy is that definition-based examples cannot consistently provide useful context information for all test samples. In contrast, random-based examples consistently offer each test sample domain-specific information relevant to the target distribution.
Comparative Analysis of Inference Time Metrics with Existing TPT Method. Table 8 reports the comparative analysis of inference time metrics with TPT [32]. All experiments are conducted on one 2080 Ti GPU, and inference time (Infer. Time) is calculated in minutes. The table shows that TPT [32] needs significant inference time due to augmenting images by 64 times. In contrast, our InCPL only uses few in-context examples (i.e., 5) without any augmentation, requiring less inference time.
@@ 5 Concusion
Limitation. We develop a novel in-context evaluation paradigm that improves the performance of existing prompt learning algorithms in a ChatGPT-like manner. Under this paradigm, the in-context evaluator operates on test samples at test-time with unaltered model parameters, where the model only processes unlabeled test sample and in-context examples served as context information, without any other samples from the same category. The proposed InCPL requires a task-specific context dataset for each individual task. This dataset creation process inherently introduces certain priors, which can potentially limit the broad applicability of our proposed methods. In future work, we will consider developing unique context datasets tailored to different tasks. This approach would enable the adaptive selection of image-label pairs to serve as in-context prompt information for various tasks.